OpenStudy (anonymous):

the sum of the angles measures of a polygon with s sides is 2340. find s

OpenStudy (anonymous):

how many sides does the polygon have?

OpenStudy (sleepyhead314):

the sum of the angles of a polygon = 180(s-2) when s = number of sides so you have 2340 = 180(s-2) do you know how to solve from there? :)

OpenStudy (jmark):

Sum of the interior angle of polygon is: 180 (s-2) = 2340 180s – 360 = 2340 180s = 2340+360 180s= 2700 s= 2700/180 s = 15 ICSE class 7 maths question paper - http://icse.edurite.com/icse-question-papers/icse-class-7-maths-question-papers.html